OpenStudy (anonymous):

At a charity dinner, money was raised by selling draw tickets for $100 each. There were 290 tickets sold. Some of the money was then given away as nine draw prizes of $1000 each. Write and Integer statment for each of the following scenarios, and solve At a charity dinner, money was raised by selling draw tickets for $100 each. There were 290 tickets sold. Some of the money was then given away as nine draw prizes of $1000 each. Write and Integer statment for each of the following scenarios, and solve @Mathematics

OpenStudy (anonymous):

100x290-(1000x9)=money made 20000$

OpenStudy (anonymous):

thankyou!, then it says: How much money remained for the charity after the prizes were awarded?

OpenStudy (anonymous):

20000$

OpenStudy (anonymous):

total money from charity event without giving away the money is 29000$

OpenStudy (anonymous):

no after the prizes were awarded

OpenStudy (anonymous):

20000$

OpenStudy (anonymous):

and why?

OpenStudy (anonymous):

100x290-(1000x9)=money made

OpenStudy (anonymous):

1000x9=prize money

OpenStudy (anonymous):

100x290=total money from tickets

OpenStudy (anonymous):

29000-9000=20000

OpenStudy (anonymous):

thx! next ques: At keeley's table there were seven other guests. Each person said they would share their winnings minus the prize ticket equally with everyone else at the table. keeley won one of the $1000 draws. How much did each person receive?

OpenStudy (anonymous):

2000-(8x100) =1200/8 =150$

OpenStudy (anonymous):

were did the 2000 come from?

OpenStudy (anonymous):

whops

OpenStudy (anonymous):

1000-(8x100) =200/8 =25$

OpenStudy (anonymous):

and the 8x100?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

tats teh amount of money teh ticket costs

OpenStudy (anonymous):

ohh ok thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

then keeley;s dad bought a table of tickets for his emplyees (seven tickets, plus one for himself.) two emplees tickets won how much money did the table give to the charity? explain.

OpenStudy (anonymous):

(7+1) x 100= money gave to charity 2x1000=money won 800-2000=money gave to charity -1200$ was given to charity

OpenStudy (anonymous):

thankyou!